OpenStudy (josee):

Determine the absolute maximum value of f(x)= xsqrt(1-x^2)-3

OpenStudy (josee):

on [-1,1]

OpenStudy (anonymous):

\[f(x)=x \sqrt{1-x^2}-3\] first we note the function is continuous on [-1,1]. Then we find the critical numbers: \[f'(x)=\sqrt{1-x^2}-\frac{x^2}{\sqrt{1-x^2}}\]The critical numbers are the values of the function where the derivative equals zero. Thus,\[2x^2=1\]And,\[x=\pm1/\sqrt{2}\]Now we check the value of the function at the end points x=-1 and x=1:\[f(1)=-3; f(-1)=-3\]The largest of -3,1/sqrt(2) and -1/sqrt(2) is your global maximum. The absolute max is \[\frac{+1}{\sqrt{2}}\]

OpenStudy (anonymous):

whoops. the value of the absolute max occures at 1/sqrt2. So plug in x=1/sqrt2 into f(x) to get the max value. I get f(x)=-5/2

OpenStudy (josee):

omggg hahaha i got up all the way to 1/sqrt2 i didnt know you had to do that last step

OpenStudy (anonymous):

lol yeah I almost forgot too. Make sure you evaluate -3 and -1/sqrt2 as well to see which is biggest/smallest. I skipped that step above too, but the solution is correct. f(1/sqrt2)=-5/2 is the max.

OpenStudy (josee):

Can you elp me with a step on this next problem? i dont wanna now the answer im just getting stuck and need a push

OpenStudy (josee):

know*

OpenStudy (anonymous):

yeah sure

OpenStudy (josee):

its find all the critical points of the function f(x)= 2sin(x)-|x| on the interval (-pi,pi) ok so i got the derivative and i set it equal to 0 right?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

but you have to deal with your absolute value of x term

OpenStudy (anonymous):

so, i would split this one into two problems: g(x)=2sinx-x on [0,pi) and h(x)=2sinx+x on (-pi,0)

OpenStudy (josee):

ok... i just took the derivative of the |x| which is x/sqrtx^2.. and when i set it equal to 0, x= pi/3

OpenStudy (anonymous):

No. The derivative of abs(x) is wrong.

OpenStudy (josee):

ohh it is? :/ damn

OpenStudy (anonymous):

If \[x \ge0\]Then\[\left| x \right|=x\]

OpenStudy (anonymous):

If x<0 then\[\left| x \right|=-x\]

OpenStudy (anonymous):

So your derivative is either +1 or -1 depending on whether x is greater than (or equal to) zero or not

OpenStudy (anonymous):

So we have two functions here, since x takes on both positive and negative values: g(x)=2sinx-x [0,pi) h(x)2sinx+x (-pi,0)

OpenStudy (josee):

so take the derivative of each function ad=nd then set that = to 0 and just combine the two?

OpenStudy (anonymous):

Taking the derivatives and setting equal to zero: On the interval [0,pi) 0=2cosx-1 cosx=1/2 x=pi/3 It's in the interval so we're good. and for the interval (-pi,0) 0=2cosx+1 cosx=-1/2 x=-2pi/3

OpenStudy (anonymous):

Your critical numbers should be \[x=\pi/3, -2\pi/3\]

OpenStudy (josee):

and 0! :D

OpenStudy (josee):

and 0! :D

OpenStudy (anonymous):

x=0 is not a critical value of this one

OpenStudy (josee):

yes it is

OpenStudy (anonymous):

Because the derivatives don't equal zero when x=0

OpenStudy (josee):

the answer was pi/3, -2pi/3 and 0 :/

OpenStudy (anonymous):

was the question asking just for the critical values?

OpenStudy (josee):

yea. find all critical values on the interval (-pi,pi)

OpenStudy (anonymous):

Oh crap! I know why

OpenStudy (anonymous):

A critical point of a function of a real variable is any value in the domain where either the function is not differentiable or its derivative is 0. The function actually isn't differentiable at x=0 because the absolute value function isn't differentiable here . You're correct x=0 too! :)

OpenStudy (anonymous):

Forgot about that part of the definition

OpenStudy (anonymous):

have you seen why abs(x) have no derivative at zero?

OpenStudy (josee):

yes :)!

OpenStudy (anonymous):

cool :))

OpenStudy (anonymous):

thanks for reminding me how to do this stuff lol

OpenStudy (josee):

thanks for helping me with my steps! and teaching me more than my professor:D